the taxes on a piece of property valued at $4,000 were $66. At this same rate, what would be the taxes on property valued at $5,500
Answer by checkley75(3666)

66/4000=X/5500
4000X=66*5500
4000X=363000
X=363000/4000
X=$90.75 IS THE TAXES ON THE 5500 VALED PROPERTY.